# JudgeAgent: Knowledge-wise and Dynamic LLM Evaluation with Agent-as-Interviewer

### v3

@@ Abstract
Current evaluation paradigms for large language models (LLMs) suffer from overestimated or biased evaluations and mismatched question difficulty, leading to incomplete evaluations of knowledge and capability boundaries, which hinder their effective application and optimization. To address these challenges, we propose Agent-as-Interviewer, a dynamic evaluation paradigm that employs LLM agents to conduct multi-turn interactions for evaluation. Unlike current benchmarking or dynamic interaction paradigms, Agent-as-Interviewer utilizes agents to invoke knowledge tools for wider and deeper knowledge in the dynamic multi-turn question generation, achieving more comprehensive evaluations of LLM’s knowledge boundaries. It also leverages agents to plan query strategies for adjustment of the question difficulty levels, enhancing the difficulty control to match the actual capabilities of target LLMs. Based on this paradigm, we develop JudgeAgent, a knowledge-wise dynamic evaluation framework that employs knowledge-driven synthesis as the agent’s tool and uses difficulty scoring as strategy guidance, thereby finally providing valuable suggestions to help targets optimize themselves. Extensive experiments validate the effectiveness of JudgeAgent’s suggestions, demonstrating that Agent-as-Interviewer can accurately identify the knowledge and capability boundaries of target models. The source code is available on https://github.com/DataArcTech/JudgeAgent.
@@ 1 Introduction
Evaluating large language models (LLMs) to understand the boundaries of their knowledge and capabilities is a critical step for their successful application across various domains (Tang et al., 2024; Yuan et al., 2023; Shi, 2024; Wang et al., 2025). Current mainstream evaluation methods for LLMs rely on static benchmark evaluations (Clark et al., 2018; Hendrycks et al., 2021; Huang et al., 2023; Lin et al., 2022; Cobbe et al., 2021; Tang & Yang, 2024), where predefined questions are posed to the target LLM and the target’s performance is evaluated by humans (Chang et al., 2024) or LLM-as-a-judge (Liu et al., 2023; Wang et al., 2023b; Zheng et al., 2023). Benefiting from the controlled question quality and straightforward workflows of the static benchmarking paradigm, developers can rapidly acquire a fundamental understanding of the strengths and limitations of various evolving LLMs, thereby facilitating swift iteration in LLM applications.
However, this benchmarking paradigm faces the bottleneck that benchmarks have become saturated more and more quickly in recent years. It took until 2023 for the accuracy record of MMLU(Hendrycks et al., 2020), which was released in 2020, to reach 80%. While the GPQA benchmark(Rein et al., 2023), published in late 2023, achieved an 80% record just within one year. The bottleneck is due to the static nature(Gu et al., 2024; Ko et al., 2024) of this paradigm. Firstly, the static paradigm restricts evaluations within a predefined knowledge scope, making it difficult to accurately measure the model’s knowledge mastery (Wang et al., 2023a; Kwan et al., 2024), leading to out-of-domain risks in practical applications. Secondly, the lack of dynamic updates increases the risk of data contamination(Schaeffer, 2023; Oren et al., 2023), where target models artificially inflate their benchmark performance by being exposed to test questions during training, leading to an overestimation of their actual capabilities. These shortcomings result in LLMs being misapplied in scenarios beyond their actual capabilities, causing the waste of resources.
Consequently, researchers begin to focus on dynamic evaluation paradigms that leverage LLMs to dynamically modify questions. Initial efforts involve using LLMs to modify questions from static benchmarks to mildly mitigate data contamination (Bai et al., 2023; Shi et al., 2025). Beyond these methods, researchers proposed paradigms that dynamically adjust questions during multi-turn interactions based on the model’s responses, enabling a deeper evaluation (Wang et al., 2023a; Kim et al., 2025). While these dynamic approaches partially address challenges in static paradigms, such as data contamination and plain evaluations, they still suffer from biased evaluation and mismatched difficulty due to simplified feedback mechanisms and knowledge limitations of evaluator LLMs, leading to a misalignment between questions and the actual range of the target’s capability. This misalignment hinders precise evaluations of the target LLM’s knowledge and capability boundaries, thereby impeding effective guidance for both application and subsequent optimization of LLMs.
Figure 1: The difference between Agent-as-Interviewer and current evaluation paradigms.
To address these challenges, we propose Agent-as-Interviewer, a dynamic evaluation paradigm that employs LLM agents to conduct multi-turn interactions for evaluation. Unlike current benchmarking or dynamic interaction paradigms, in the dynamic follow-up question generation based on the target’s responses in multi-turn interactions, Agent-as-Interviewer utilizes agents to call knowledge tools for wider and deeper knowledge, achieving more complete evaluations of the target LLM’s knowledge and capability boundaries. It also leverages agents to plan query strategies for adjustment of the question difficulty levels, enhancing the difficulty control to match the actual capabilities of target LLMs. Based on this paradigm, we develop JudgeAgent, a knowledge-wise dynamic evaluation framework that employs knowledge-driven synthesis as the agent’s tool and uses difficulty scoring as strategy guidance in multi-turn dynamic interactions, thereby finally providing valuable suggestions to help targets optimize themselves.
@@
  • Agent-as-Interviewer paradigm: We propose a dynamic evaluation paradigm that employs LLM agents to conduct multi-turn interactions for evaluation, addressing the challenges of incomplete evaluation and inadequate difficulty control in current paradigms.
  • Knowledge-wise evaluation framework: We introduce JudgeAgent, a knowledge-wise dynamic evaluation framework based on Agent-as-Interviewer, to provide valuable suggestions and precise guidance for the optimization of target LLMs.
  • Validation of the paradigm: We conduct thorough experiments and analysis to validate the effectiveness of Agent-as-Interviewer and JudgeAgent’s suggestions.
@@ 2.1 Static Benchmark-based Evaluation
These methods employ pre-constructed benchmarks to evaluate LLMs in specific tasks, using formats such as multiple-choice, question-answer(Q&A), or prompts for performing tasks. For multiple-choice (Clark et al., 2018; Hendrycks et al., 2021; Huang et al., 2023) or Q&A (Lin et al., 2022; Cobbe et al., 2021; Tang & Yang, 2024) formats, the benchmark provides correct answers and measures the target’s capability by its accuracy. For task-execution format, the benchmark measures the model by the correctness and plausibility assessed by human (Chang et al., 2024) or LLM-as-a-judge (Wang et al., 2023b; Zheng et al., 2023) based on reference answers. These methods ensure controllable question quality and rapid pipelines, but they are also susceptible to data contamination, which undermines the validity of evaluations.
@@ 2.2 Dynamic LLM-based Evaluation
These methods leverage LLMs to dynamically generate evaluation questions. Bai et al. (2023) employs LLMs as examiners to generate questions based on Google Trends categories. SafetyQuizzer(Shi et al., 2025) formulates questions with current events retrieved from search engines to maximize the timeliness. These methods focus on the knowledge sources, but lack interaction with target models. Therefore, researchers attempt to generate follow-up questions based on responses in multi-round Q&A sessions. KIEval (Yu et al., 2024) and LLM-as-an-Interviewer (Kim et al., 2025) leverage LLMs to generate follow-up questions and provide feedback based on the target’s responses in multi-turn interactions, thereby probing the depth of the target’s knowledge.
Compared to static methods, their advantage is that target models cannot prepare for generated questions, thereby producing more genuine evaluations. However, these dynamic processes are often centered on a single question, and all the generations rely entirely on the LLM, resulting in a limited scope of knowledge coverage and pool quality control. Moreover, these methods focus on the content of questions, without properly calibrating question difficulty to the target’s capability range, leading to biased evaluations. To address these challenges, we propose Agent-as-Interviewer and develop JudgeAgent based on this paradigm, which performs fine-grained evaluation through knowledge-wise generation and adaptive difficulty adjustment in dynamic multi-turn interactions.
@@ 3 Methodology
We introduce Agent-as-Interviewer, which utilizes LLM agents to simulate the entire interview for the dynamic evaluation of LLMs, and develop JudgeAgent based on this paradigm. In this section, we use JudgeAgent as an example to illustrate the workflow of this paradigm. The workflow comprises three core components: (1) Benchmark Grading: Fundamentally estimating the target’s capability range through testing on public static benchmarks. (2) Interactive Extension: Using knowledge-driven data synthesis to dynamically generate follow-up questions and update the difficulty levels based on the target’s capability estimations. (3) Evaluation Feedback: Evaluating the target’s deficiencies in multiple dimensions and providing actionable suggestions for optimization. The overall framework is illustrated in Figure 2, and detailed as pseudo code in Algorithm 1.
@@ 3.1 Benchmark Grading
Analogous to a written test before an interview, JudgeAgent forms a base understanding of the target model’s capabilities by evaluating its performance on a publicly available static benchmark.
JudgeAgent first partitions the benchmark questions into batches and then prompts the target to answer the questions batch by batch. After assessing the target’s answers, JudgeAgent estimates the target’s capability based on its performance in each batch. Within each batch, each correct answer improves the total score, and the average score determines the follow-up question’s difficulty in the extension stage, including Easy, Medium, and Hard. Let $d_{1},d_{2},d_{3}]$ respectively denote the difficulty Easy, Medium, and Hard, $n_{it},n_{if}$ respectively represent the number of correct and incorrect answers for $d_{i}$-level questions, $c_{it}$ be the score gain for a correct answer at $d_{i}$, and $t_{ij}$ be the score threshold between $d_{i}$ and $d_{j}$. The difficulty control mechanism is formalized as follows:
[…]
In JudgeAgent, we set $c_{1t},c_{2t},c_{3t}]=[1,1.5,2]$ and $t_{12},t_{23}]=[0.5,1]$. For benchmarks without difficulty labels, the default score gain is $c_{2t}$. The mechanism’s detailed design rules, analysis, and proofs are available in Appendix B.
@@ 3.2 Interactive Extension
This process is an ”interview” where JudgeAgent conducts an in-depth evaluation of target models. Specifically, JudgeAgent iteratively generates questions based on the target’s responses and dynamically controls the difficulty. The process consists of three steps: Relevant Knowledge Retrieval, Difficulty-Adaptive Question Generation, and Capability Estimation.
@@
Relevant Knowledge Retrieval: Based on seed question batches, JudgeAgent retrieves related knowledge from the benchmark’s knowledge base. If only reference texts of seed questions are used as background knowledge for generation, the generated questions will be highly similar to the original questions, with limited knowledge scopes. Inspired by SoG(Jiang et al., 2025), JudgeAgent employs a context-graph–based sampling approach to get background texts. The context graph is constructed from the benchmark’s knowledge base (e.g., the collection of all reference texts) by segmenting texts into chunks and extracting entities as nodes, with entities in the same text as neighbors. The detailed construction is illustrated in Appendix C and Algorithm 3.
@@
During sampling, JudgeAgent first extracts entities from seed questions and finds the most similar entity $e$ on the graph. In each hop of sampling, JudgeAgent retrieves the most relevant chunks of $e$ to the question, randomly selects a chunk $c$, treats $(e,c)$ as the root, and $e$ is used as the entity for the next hop. This process is repeated for $N$ hops to form a knowledge path with breadth and depth, and the chunks along the path are concatenated to form the background knowledge. The context graph ensures both knowledge breadth and depth, and the greedy similarity strategy ensures relevance to seed questions. Entity extraction is performed by an LLM, which is GPT-4.1 in our experiments, with prompts detailed in Appendix G. This step provides the appropriate knowledge background for JudgeAgent to generate knowledge-wise questions.
@@
Difficulty-Adaptive Question Generation: Unlike other dynamic methods, JudgeAgent incorporates the target’s capability estimation into question generation to enable more fine-grained evaluations. A fixed-difficulty test can identify underperforming models, but can hardly comprehensively evaluate a model’s knowledge levels. For instance, when an overly challenging test evaluates a model with limited knowledge, the evaluation may reveal that the model performs poorly, but fails to delineate the actual range of its knowledge. Therefore, JudgeAgent dynamically adjusts the target’s capability estimation and the follow-up question’s difficulty based on Eq 1. This adaptive approach allows for a precise and refined understanding of the target’s knowledge boundaries.
@@
Based on the capability estimation, JudgeAgent adjusts the question difficulty as Easy, Medium, or Hard, and explicitly specifies each difficulty’s requirements in the prompt. For Easy, the focus is on assessing information retrieval skills, primarily through cloze-style questions based on the original text, like ”Which virus primarily causes HFMD?”. For Medium, the emphasis is on understanding key concepts, with questions involving basic inference, like ”Which description is correct regarding meiosis II?” For Hard, the questions are designed to encourage deep thinking and complex logical analysis, like ”A patient ingested a toxic substance… The doctor employed… which type of treatment is likely lacking?”. Detailed prompt designs are provided in Appendix G.
Capability Estimation: In each round, JudgeAgent presents the generated questions to the target model and collects its responses. Based on all responses in the above two stages, JudgeAgent computes scores according to the scoring scheme described in Section 3.1, considering both difficulty and correctness. The average score is used to determine the difficulty for the next round’s generation.

### v4

# JudgeAgent: Beyond Static Benchmarks for Knowledge-Driven and Dynamic LLM Evaluation

@@ Abstract
Current evaluation methods for large language models (LLMs) primarily rely on static benchmarks, presenting two major challenges: limited knowledge coverage and fixed difficulties that mismatch with the evaluated LLMs. These limitations lead to superficial assessments of LLM knowledge, thereby impeding the targeted model optimizations. To bridge this gap, we propose JudgeAgent, a knowledge-driven and dynamic evaluation framework for LLMs. To address the challenge of limited knowledge coverage, JudgeAgent leverages LLM agents equipped with context graphs to traverse knowledge structures systematically for question generation. Furthermore, to mitigate data contamination and difficulty mismatch, it adopts a difficulty-adaptive and multi-turn interview mechanism. Thereby, JudgeAgent can achieve comprehensive evaluations and facilitate more effective improvement of LLMs. Empirical results demonstrate that JudgeAgent enables more comprehensive evaluations and facilitates effective model iterations, highlighting the potential of this knowledge-driven and dynamic evaluation paradigm. The source code is available on https://github.com/DataArcTech/JudgeAgent.
JudgeAgent: Beyond Static Benchmarks for Knowledge-Driven and Dynamic LLM Evaluation
Zhichao Shi1,2,3,4thanks: Both authors contributed equally to this research., Xuhui Jiang1,211footnotemark: 1, Chengjin Xu1,2, Cangli Yao1,2, Shengjia Ma1,2, Yinghan Shen4, Zixuan Li4, Jian Guo2, Yuanzhuo Wang4thanks: Corresponding author, 1DataArc Tech Ltd. 2IDEA Research, International Digital Economy Academy 3School of Advanced Interdisciplinary Sciences, UCAS 4State Key Lab of AI Safety, Institute of Computing Technology, CAS
@@ 1 Introduction
Evaluating Large Language Models (LLMs) to verify their knowledge is a crucial step for their successful application across various domains.(Tang et al., 2024; Yuan et al., 2023; Shi, 2024; Wang et al., 2025). Current evaluation methods mainly rely on static benchmarks(Clark et al., 2018; Hendrycks et al., 2021; Huang et al., 2023; Lin et al., 2022; Cobbe et al., 2021; Tang and Yang, 2024), where LLMs are evaluated by their performance on answering predefined static questions. In the early stages, benefiting from the controllable question quality and rapid workflows of static benchmarking, developers can quickly grasp the core competencies of evolving LLMs, thereby accelerating iteration cycles. However, due to the static nature(Gu et al., 2024; Ko et al., 2024), these benchmarks have become saturated more and more quickly in recent years. For example, it took 3 years for MMLU(Hendrycks et al., 2020) to reach 80% accuracy record, while GPQA(Rein et al., 2023) achieved the same record just within one year. The static nature restricts the evaluations within a predefined and limited knowledge scope(Wang et al., 2023a; Kwan et al., 2024), and increases the risk of data contamination(Schaeffer, 2023; Oren et al., 2023), resulting in LLMs being misapplied in scenarios out of their knowledge.
Figure 1: The difference between JudgeAgent and current evaluation paradigms.
Consequently, researchers attempted to modify static benchmarks using LLMs(Bai et al., 2023; Shi et al., 2025). However, they still face two major challenges. (1) Limited knowledge coverage: Constrained by the evaluator LLM’s knowledge limitations and inherent randomness, the knowledge extended by the LLM from static benchmarks remains uncontrollable and limited, making it difficult to achieve comprehensive knowledge evaluations. (2) Mismatch between question difficulty and the evaluated LLMs: Current mainstream methods lack adaptive adjustments to question difficulty. Thus, the question difficulty may significantly deviate from the actual knowledge mastery of the evaluated LLMs, making the evaluations harder to precisely reflect the model’s capabilities. Thus, a new knowledge-driven and dynamic evaluation framework is needed.
In this paper, we propose JudgeAgent, a knowledge-driven and dynamic evaluation framework for LLMs. To address the challenge of limited knowledge coverage, JudgeAgent leverages LLM agents with context graphs to traverse knowledge structures at greater breadth and depth. Through graph sampling strategies and knowledge-driven data synthesis, JudgeAgent generates questions with broader and deeper knowledge coverage, further exploring potential knowledge deficiencies in evaluated LLMs. Furthermore, to mitigate data contamination and difficulty mismatch, JudgeAgent dynamically adjusts the capability estimation and generates difficulty-adaptive questions based on the target’s responses in multi-turn interviews. Thereby, JudgeAgent can provide comprehensive evaluations and facilitate more effective subsequent improvement of evaluated LLMs.
Extensive experiments and analysis validate that JudgeAgent can provide effective and precise evaluation results. Our in-depth analysis also reveals the significant potential of this evaluation paradigm. It offers valuable insights into how additional knowledge-driven and dynamic strategies for evaluator LLMs can be designed, which enhance the quality of dynamic evaluation results.
@@
  • We introduce JudgeAgent, a knowledge-driven and dynamic evaluation framework for LLMs. JudgeAgent leverages LLM agents with context graphs to traverse knowledge structures and perform dynamic multi-turn interviews, achieving comprehensive knowledge evaluations.
  • We propose a difficulty-adaptive and multi-turn evaluation mechanism that simulates an interactive interview, enabling the evaluations to more precisely reflect the knowledge mastery of the evaluated LLM.
  • We conduct extensive experiments to validate the effectiveness of JudgeAgent and highlight its significant potential for the dynamic evaluation paradigm.
@@ 2.1 Static Benchmark-based Evaluation
These methods employ pre-constructed benchmarks to evaluate LLMs, using formats such as multiple-choice, question-answer(Q&A), or prompts for performing tasks. For multiple-choice (Clark et al., 2018; Hendrycks et al., 2021; Huang et al., 2023) or Q&A (Lin et al., 2022; Cobbe et al., 2021; Tang and Yang, 2024), the benchmark provides correct answers and assesses the target by the accuracy. For task-execution format, the benchmark measures the model by the metrics assessed by human (Chang et al., 2024) or LLM-as-a-judge (Wang et al., 2023b; Zheng et al., 2023). These methods ensure controllable question quality, but they are also susceptible to data contamination, which undermines the validity of evaluations.
@@ 2.2 Dynamic LLM-based Evaluation
With the deepening application of LLMs, in addition to mainstream static methods, recent researchers have begun to leverage LLMs to dynamically generate evaluation questions. Bai et al. (2023) employs LLMs as examiners to generate questions based on Google Trends categories. SafetyQuizzer(Shi et al., 2025) formulates questions with current events from search engines to maximize the timeliness. KIEval(Yu et al., 2024), and LLM-as-an-Interviewer (Kim et al., 2025) leverage LLMs to generate follow-up questions based on the target’s responses in multi-turn interactions.
Compared to static methods, these dynamic methods address the challenge of data contamination and provide more genuine evaluations. However, the generations of these methods rely mainly on the LLM, resulting in limited knowledge coverage and pool quality control. Moreover, these methods lack adaptive adjustments of question difficulty, resulting in a deviation between the evaluations and the actual capabilities of target LLMs. This mismatch leads to biased subsequent optimization efforts and reduced efficiency.
Thus, we propose a knowledge-driven and dynamic evaluation framework named JudgeAgent. To expand knowledge coverage, JudgeAgent leverages LLM agents with context graphs to traverse knowledge structures and generate questions. To mitigate data contamination and difficulty mismatch, JudgeAgent adopts a difficulty-adaptive multi-turn interview mechanism to conduct dynamic evaluations. Thus, JudgeAgent provides comprehensive evaluations and facilitates subsequent improvement of LLMs.
@@ 3 Methodology
We introduce JudgeAgent, a knowledge-driven dynamic evaluation framework, which utilizes LLM agents with context graphs to simulate the entire interview of target LLMs. The workflow comprises three core components: (1) Benchmark Grading: Fundamentally assessing target LLMs through testing on public static benchmarks. (2) Interactive Extension: Using knowledge-driven synthesis to dynamically generate follow-up questions and update the difficulty based on the estimations of target LLMs. (3) Evaluation Feedback: Evaluating the target’s deficiencies in multiple dimensions and providing actionable suggestions. The overall framework is illustrated in Figure 2, and detailed as pseudo code in Algorithm 1.
@@ 3.1 Benchmark Grading
To address the challenge of difficulty mismatch in prior dynamic evaluation, JudgeAgent first evaluates the target LLM on public static benchmarks and obtains a base score through a difficulty scoring mechanism. Similar to a written test before an interview, this allows JudgeAgent to form a basic capability estimation of the target LLM.
JudgeAgent first partitions the benchmark questions into batches. Then it prompts the target LLM to answer the questions batch by batch. After assessing the responses, JudgeAgent estimates the target’s capability based on its performance in each batch. Inspired by the common real-world practice of scoring student exams, we design a linear difficulty control mechanism. Within each batch, each correct answer improves the total score, and the average score determines the follow-up question’s difficulty. Let $d_{1},d_{2},d_{3}]$ respectively denote the difficulty Easy, Medium, and Hard, for the score gain $c_{it}$ for a correct answer at different difficulty levels $d_{i}$, we set $c_{1t},c_{2t},c_{3t}]=[1,1.5,2]$. And for the threshold $t_{ij}$ between $d_{i}$ and $d_{j}$, we set $t_{12},t_{23}]=[0.5,1]$. Once the dynamic average score exceeds the threshold $t_{ij}$, the difficulty of the next generated question will be set to $d_{j}$. The mechanism’s detailed design, analysis, and proofs are available in Appendix B.
@@ 3.2 Interactive Extension
This process is an "interview" where JudgeAgent conducts a knowledge-driven and difficulty-adaptive multi-turn evaluation. Specifically, JudgeAgent dynamically adjusts the difficulty and generates questions with expanded knowledge based on the responses. Furthermore, JudgeAgent explores the potential knowledge gaps of target LLMs through a multi-turn interaction process. Each iteration in this process consists of three steps: Relevant Knowledge Retrieval, Difficulty-Adaptive Question Generation, and Capability Estimation.
@@
Relevant Knowledge Retrieval: To achieve knowledge-driven evaluation, JudgeAgent invokes external knowledge and constructs context graphs to facilitate the association and extension of knowledge during evaluation. If only reference texts of seed questions are used for generation, the generated questions will be highly similar to the original ones. Inspired by SoG(Jiang et al., 2025), JudgeAgent employs a context-graph–based sampling approach to get background texts. The context graph is constructed from the benchmark’s knowledge base with entities and chunks extracted from texts as nodes. The entities in the same text will be linked on the graph. The detailed construction is illustrated in Appendix C and Algorithm 3.
@@
During sampling, JudgeAgent first extracts entities from seed questions and finds the most similar entity $e$ on the graph. In each hop, JudgeAgent retrieves the most relevant chunks of $e$ to the question, and randomly selects a chunk $c$. $e$ is the seed entity of the next hop, and $c$ is added to the path. After $N$ hops, we sample a path with breadth and depth, and the chunks on the path are concatenated to form the knowledge for generation. The context graph ensures both knowledge breadth and depth, and the greedy similarity strategy ensures relevance to seed questions. Entity extraction is performed by GPT-4.1, with prompts detailed in Appendix G.
@@
Difficulty-Adaptive Question Generation: To mitigate data contamination and difficulty mismatch, JudgeAgent conducts multi-turn interactive interviews to explore potential deficiencies and incorporates adaptive difficulty adjustments into question generation. A fixed-difficulty test can identify underperforming models, but can hardly comprehensively evaluate a model’s actual knowledge mastery. For instance, when an overly challenging test evaluates a model with limited knowledge, the evaluation may reveal that the model performs poorly, but fails to delineate the actual range of its knowledge. Therefore, JudgeAgent dynamically adjusts the follow-up question’s difficulty based on capability estimation through the difficulty control mechanism in Section 3.1.
@@
Based on the capability estimation, JudgeAgent adjusts the question difficulty as Easy, Medium, or Hard, and explicitly specifies each difficulty’s requirements in the prompt. For Easy, the focus is on assessing information retrieval skills, primarily through cloze-style questions based on the original text, like "Which virus primarily causes HFMD?". For Medium, the emphasis is on understanding key concepts, with questions involving basic inference, like "Which description is correct regarding meiosis II?" For Hard, the questions are designed to encourage deep thinking and complex logical analysis, like "A patient ingested a toxic substance… The doctor employed… which type of treatment is likely lacking?". Detailed prompt designs are provided in Appendix G, and the validation of question quality is described and analyzed in Appendix E.4.
Capability Estimation: JudgeAgent queries the target LLM with generated questions. Based on LLM’s responses, JudgeAgent computes average difficulty scores with the strategy in Section 3.1, and determines the difficulty for the next round’s generation. Through this adaptive multi-turn mechanism, JudgeAgent progressively aligns the assessment with the true capabilities of the target LLM, thereby achieving more precise evaluations.